# Modules as effective nodes in coarse-grained networks of Kuramoto oscillators

### v1

@@ VI Conclusions
Network properties like size, topology and connectivity are important to explain the disparity in synchronization between real and coarse-grained networks. Zachary?s network e.g. has higher connectivity to number of nodes ratio than C. Elegans network. It has also been shown that to retain synchronization, some Laplacian eigenvalues should be preserved between the original and coarse-grained networks [49]. Understanding the role of these properties in the coarse-graining process will help to unravel under what circumstances treating groups of nodes as single oscillators is a valid assumption.
@@ Acknowledgements.
It is a pleasure to thank J. A. Brum and Guilherme F. Arruda for helpful suggestions. This work was partially supported by FAPESP, grant 2021/14335-0 (MAMA) CAPES, grant 88887.706116/2022-00 (LLB), and CNPq, grant 303814/2023-3 (MAMA).

### v2

@@ VII Discussion
We have seen that network properties like topology and link density are important for the proper application of the coarse graining procedure. Modules that are loosely connected do not synchronize and might affect the accuracy of the results. In the examples shown here for the C. Elegans, such modules were small and did not compromise the quality of the results. It has been shown, for example, that to retain synchronization, some Laplacian eigenvalues should be preserved between the original and coarse-grained networks [59]. Understanding the role of these properties in the coarse-graining process will help to unravel under what circumstances treating groups of nodes as single oscillators is a valid assumption.
@@ Acknowledgements.
It is a pleasure to thank J. A. Brum and Guilherme F. Arruda for helpful suggestions. Relevant codes and documentation are available at open access GitHub repository [60]. This work was partially supported by FAPESP, grant 2021/14335-0 (MAMA) CAPES, grant 88887.706116/2022-00, and CNPq, grant 303814/2023-3 (MAMA).